# The Host Galaxies of Radio AGN: New Views from Combining LoTSS and MaNGA Observations

### v1

@@ 5.2 Disparity between the jet age and quenching time
Most of the RDAGN in this study lack sufficient frequency coverage and spatial information to accurately determine their spectral and dynamical jet ages. However, even for extremely old jets with well-calibrated dynamical and spectral ages, the jet ages are less than 1 Gyr (Turner et al. 2018), which is significantly shorter than the quiescent period observed in our RDAGN hosts. As depicted in Fig. 7, RDAGN hosts are observed to have become quenched approximately 3 Gyr ago. If both timescale measurements are accurate within factors of a few, our results suggest that the onset of radio jet activity likely occurred after the host galaxy became quenched, contradicting the scenario where AGN appear and then suppress star formation.
This then implies that the observed RDAGN activity is not responsible for the host quenching. Several recent studies have argued that different AGN populations may represent different evolution phases of AGN, and the radio AGN represent the very late stages of a cycle of AGN activity (e.g., Klindt et al. 2019). If so, this would explain why the radio jet activities may occur later than the quenching, since the star formation suppression happened in AGN’s early stages when the radiation feedback is strong. Even with the short timescale of radio jet activity, we still find high RDAGN incidence in the most massive galaxies (50%, see Fig. 5), which indicates that the radio mode AGN in massive galaxies are switching on and off rapidly with a high duty cycle to be observed. And the current outburst is only one of the many AGN activities which can date back closer to the quenching time. The timescales of AGN evolution are still poorly constrained, however.
@@ 5.3 The triggering of radio AGN
In Figure 5, we showed that the fraction of galaxies that host radio-loud AGN is a very strong function of the stellar mass of the galaxy. In this section, we attempt to pinpoint the underlying cause for why radio AGN activity is more common in massive galaxies by including additional parameters in the analysis, namely SMBH mass, the cosmic environment, and the galaxy morphology. The SMBH mass is estimated by the relation used in McConnell & Ma (2013). The is fitted from the MaNGA spectra using DAP (Westfall et al. 2019). We adopt the numbers of neighbors within the projected 200 kpc radius and line-of-sight as an approximation for the environmental impact, which are taken from the MaNGA-GEMA catalog333https://data.sdss.org/datamodel/files/MANGA_GEMA/GEMA_VER (Argudo-Fernádez et al. in prep). We use the Hubble type visually classified by Vázquez-Mata et al. (2022) and divided galaxies into elliptical, S0, and late-type (spiral and irregular galaxies).
We calculate the incidence of RDAGN as a function of stellar mass, and plot the results for subsamples divided by the 3 additional parameters in Fig.13. In the upper panel, the RDAGN incidence is clearly higher in high systems at fixed stellar mass. The middle and lower panels indicate that at fixed stellar mass, RDAGN prefer denser cosmic environment and late-type galaxies, which is consistent with previous findings (e.g., Best et al. 2005a). The trends as a function of at fixed stellar mass are by far the strongest ones, suggesting that black hole mass is an important parameter, after stellar mass, that controls the fraction of galaxies with radio mode AGN activity. We note that at a similar Eddington ratio, the more massive black holes will have higher luminosity and thus are more likely to be detected. Sabater et al. (2019) used a larger single-fiber observed sample and showed that the RDAGN incidence dpendence on is marginal. This inconsistency is partly due to that they made a radio luminosity cut on the radio AGN sample. After applying a similar cut (), our fraction values generally agree with the results in Sabater et al. (2019) ( at ).
@@ 5.4 Implications for the AGN feedback scenario
Because radio AGN are located in denser environments, at fixed mass, they are likely to have assembled earlier and faster. The observed higher stellar dispersion suggests that their central SMBHs are more massive, indicating possible stronger AGN activity in the past. This past AGN activity may have played an important role in quenching the host galaxies through gas heating, fast consumption of the gas, or gas outflows (e.g., Silk et al. 2024). The early-type morphology and low angular momentum (Zheng et al. 2023) indicates past galaxy-galaxy merger events, which may trigger the past AGN activity and also the fast mass growth, as well as the morphological transition from disk to spheroid (e.g. Rothberg & Joseph 2004).
@@ 5.5 Future directions
This work is based on the radio images and optical IFUs with spatial resolutions on the order of arcseconds. In the future, new telescopes and the update of current facilities can carry out large radio sky surveys with sub-arcsec resolution, e.g., The International LOFAR Two-metre Sky Survey444https://lofar-surveys.org/ilotss.html, the Next Generation Very Large Array (ngVLA, Di Francesco et al. 2019), the Square Kilometre Array555https://www.skao.int/ (SKA), etc.. At low redshifts, they can resolve the jet features on physical scales of a few hundred of parsecs. As discussed in Section 4.1, the higher resolution is also essential to decompose the faint radio AGN in star-forming galaxies, which can help us to understand the radio AGN luminosity function in the faint end. The wide frequency coverage of different instruments can also help us build the full synchrotron SED of radio AGN, which is essential for constraining jet ages.
Meanwhile, there are more and more deep and high-resolution IFU samples. These improvements can reveal the detailed process about how the jets from AGN interact with ISM. High-resolution IFU is also necessary to distinguish the jet-related emission and the AGN photoionization emission. In addition to advancements in radio and optical observations, future large-scale surveys at other wavelengths, particularly in the X-ray and infrared, are expected to match the resolution and sensitivity achieved at optical and radio bands. This multi-wavelength approach will provide comprehensive information about the different component structures of AGN, allowing for a more complete understanding of the structure and the evolution of AGN.
@@ 6 Summary
We build the largest optical IFU - radio continuum sample to date consisting of 5578 galaxies by matching the MaNGA survey with LoTSS-DR2 observations. We revisit the tight linear SFR - radio luminosity relation using IFU data and find an intrinsic error of dex. Using this relation, we identify 616 radio AGN with excessive radio emission, and calculate their jet luminosities. We analyze their global star formation rates and their correlation with jet luminosity. We then derive the radially resolved mass assembly history by fitting the IFU spectra in radial bins and comparing the difference between radio AGN hosts and their control samples We also stack the spectra in different radial bins to investigate the impact of radio AGN on the emission lines of their host galaxies. Our results can be summarized as follows:
@@
Stacked spectra comparisons reveal weak emission line excess exclusively in the central region of radio AGN hosts, especially in low mass RDAGN. The high ratio of the excessive [N II] and H supports the scenario that the emission line excess is linked to AGN or shock ionization, but it remains to be understood why the excess is not dependent upon jet luminosity. Low luminosity radio AGN hosts show emission line deficiencies in their outskirts, likely indicating that RDAGN have weaker star formation in those regions.
@@ Appendix A Fitting the -SFR relation
To build the -SFR relation, the key is to select a sample of pure star-forming galaxies with reliable SFR measurement and without radio jet contribution. Using the integral field spectra from MaNGA, the SFR can be derived from the H line, and corrected for dust attenuation on sub-galactic scales. There are galaxies in which the H is not dominated by star formation but instead is possibly contaminated by narrow-line AGN or diffuse ionized gas. We exclude these galaxies using the widely used emission line diagnostic (BPT diagram, Baldwin et al. 1981) based on the [N II]6585/H and [O III]5008/H line ratios measured from the central 2.5″MaNGA spaxels. Specifically, our ‘pure-SFG’ sample only include galaxies locating in the star-forming region of the BPT diagram, using the separation diagnostic derived by (Kauffmann et al. 2003) as well as showing nuclear H equivalent width larger than 3 (Cid Fernandes et al. 2011). Then we visually check the LoTSS image of these galaxies and exclude those showing jet-like features and require the signal-to-noise ratio of radio detection larger than 5. This way we get a sample of 906 secure SFGs with reliable SFR and radio luminosity measurements. Fig. 14 shows the emission line selection of this SFG sample. BPT is the distance to the division line in the BPT diagram (Equation 1 in Kauffmann et al. 2003), and the negative values represent that galaxies locate in the star-forming region. As a result of all the constrains, this SFG sample has a SFR limit about 0.1 (see Fig. 2), thus our result about this relation may not be applicable for SFGs with SFRs lower than this value.
We set three parameters in the linear fitting of the -SFR relation. In addition to the slope and constant, we introduce a intrinsic scatter parameter, which should be the scatter level of the -SFR relation after excluding the effect from measurement errors. Our fitting follows the method in Hogg et al. (2010). The Markov chain Monte Carlo (MCMC) sampling result is shown in Fig. 15. The intrinsic error is found to be small, about 0.23 dex, which indicates that the radio emission is good tracer for galaxy global SFR. The relation is listed in Equation 1. At the 3 level, our result is consistent with previous results based on different SFR tracers and radio data from different frequencies (Wang et al. 2016; Brown et al. 2017; Gürkan et al. 2018; Smith et al. 2021; Heesen et al. 2022; Best et al. 2023).

### v3

@@ 5.2 Disparity between the jet age and quenching time
Most of the RDAGN in this study lack sufficient frequency coverage and spatial information to accurately determine their spectral and dynamical jet ages. However, even for extremely old jets with well-calibrated dynamical and spectral ages, the jet ages are less than 1 Gyr (Turner et al. 2018), which is significantly shorter than the quiescent period observed in our RDAGN hosts. As depicted in Fig. 7, RDAGN hosts are observed to have become quenched approximately 5 Gyr ago. If both timescale measurements are accurate within factors of a few, our results suggest that the onset of radio jet activity likely occurred after the host galaxy became quenched, contradicting the scenario where AGN appear and then suppress star formation.
This then implies that the observed RDAGN activity is not responsible for the host quenching. Several recent studies have argued that different AGN populations may represent different evolution phases of AGN, and the radio AGN represent the very late stages of a cycle of AGN activity (e.g., Klindt et al. 2019; Albán et al. 2024). If so, this would explain why the radio jet activities may occur later than the quenching, since the star formation suppression happened in AGN’s early stages when the radiation feedback is strong. Even with the short timescale of radio jet activity, we still find high RDAGN incidence in the most massive galaxies (50%, see Fig. 5), which indicates that radio AGN are common in massive galaxies with a high duty cycle, and the radio mode might be an important ingredient in AGN activities. However, it is still difficult to constrain the duty cycle and evolution timescale of AGN according to the observations.
@@ 5.3 The triggering of radio AGN
In Figure 5, we showed that the fraction of galaxies that host radio-loud AGN is a very strong function of the stellar mass of the galaxy. In this section, we attempt to pinpoint the underlying cause for why radio AGN activity is more common in massive galaxies by including additional parameters in the analysis, namely SMBH mass, the cosmic environment, and the galaxy morphology. The SMBH mass is estimated by the relation used in McConnell & Ma (2013). The is fitted from the MaNGA spectra using DAP (Westfall et al. 2019). We adopt the numbers of neighbors within the projected 200 kpc radius and line-of-sight as an approximation for the environmental impact, which are taken from the MaNGA-GEMA catalog666https://data.sdss.org/datamodel/files/MANGA_GEMA/GEMA_VER (Argudo-Fernádez et al. in prep). We use the Hubble type visually classified by Vázquez-Mata et al. (2022) and divided galaxies into elliptical, S0, and late-type (spiral and irregular galaxies).
We calculate the incidence of RDAGN as a function of stellar mass, and plot the results for subsamples divided by the 3 additional parameters in Fig.14. In the upper panel, the RDAGN incidence is clearly higher in high systems at fixed stellar mass. The middle and lower panels indicate that at fixed stellar mass, RDAGN prefer denser cosmic environment and early-type galaxies, which is consistent with previous findings (e.g., Best et al. 2005a). The trends as a function of at fixed stellar mass are by far the strongest ones, suggesting that black hole mass is an important parameter, after stellar mass, that controls the fraction of galaxies with radio mode AGN activity. We note that at a similar Eddington ratio, the more massive black holes will have higher luminosity and thus are more likely to be detected. Sabater et al. (2019) used a larger single-fiber observed sample and showed that the RDAGN incidence dependence on is marginal. This inconsistency is partly due to that they made a radio luminosity cut on the radio AGN sample. After applying a similar cut (), our fraction values generally agree with the results in Sabater et al. (2019) ( at ). However, the dependence of the RDAGN incidence on still exists. Possible reasons could be our much smaller sample size, and/or the different redshift range ( versus ).
@@ 5.4 Implications for the AGN feedback scenario
Because radio AGN are located in denser environments, at fixed mass, they are likely to have assembled earlier and faster. The observed higher stellar dispersion suggests that their central SMBHs are more massive, indicating possible stronger AGN activity in the past. This past AGN activity may have played an important role in quenching the host galaxies through gas heating, fast consumption of the gas, or gas outflows (e.g., Silk et al. 2024). The early-type morphology and low angular momentum (Zheng et al. 2023) indicates past galaxy-galaxy merger events, which may trigger the past AGN activity and also the fast mass growth, as well as the morphological transition from disk to spheroid (e.g. Rothberg & Joseph 2004). This is also consistent with some simulation results which suggest that episodic AGN events, rather than one time AGN event, contribute to the final quenching of the host galaxies (e.g. Piotrowska et al. 2022; Harrison & Ramos Almeida 2024).
@@ 5.5 Future directions
This work is based on the radio images and optical IFUs with spatial resolutions on the order of arcseconds. In the future, new telescopes and the update of current facilities can carry out large radio sky surveys with sub-arcsec resolution, e.g., The International LOFAR Two-metre Sky Survey777https://lofar-surveys.org/ilotss.html, the Next Generation Very Large Array (ngVLA, Di Francesco et al. 2019), the Square Kilometre Array888https://www.skao.int/ (SKA), etc.. At low redshifts, they can resolve the jet features on physical scales of a few hundred of parsecs. As discussed in Section 4.1, the higher resolution is also essential to decompose the faint radio population in AGN and star-forming galaxies. This can help us to understand the radio AGN luminosity function in the faint end. The wide frequency coverage of different instruments can also help us build the full synchrotron SED of radio AGN, which is essential for constraining jet ages.
Meanwhile, there are more and more deep and high-resolution IFU samples. Future high-resolution IFUs can provide emission line maps around the AGN at parsec-scale resolution. Combined with high-resolution radio images, the morphology of the radio continuum and emission lines may reveal the detailed process of how the jets from AGN interact with the ISM. In addition to advancements in radio and optical observations, future large-scale surveys at other wavelengths, particularly in the X-ray and infrared, are expected to match the resolution and sensitivity achieved at optical and radio bands. This multi-wavelength approach will provide comprehensive information about the different component structures of AGN, allowing for a more complete understanding of the structure and the evolution of AGN.
@@ 6 Summary
We build the largest optical IFU - radio continuum sample to date consisting of 5548 galaxies by matching the MaNGA survey with LoTSS-DR2 observations. We revisit the tight linear SFR - radio luminosity relation using IFU data and find an intrinsic error of dex. Using this relation, we identify 616 radio AGN with excessive radio emission, and calculate their jet luminosities. We analyze their global star formation rates and their correlation with jet luminosity. We then derive the radially resolved mass assembly history by fitting the IFU spectra in radial bins and comparing the difference between radio AGN hosts and their control samples We also stack the spectra in different radial bins to investigate the impact of radio AGN on the emission lines of their host galaxies. Our results can be summarized as follows:
@@
Stacked spectra comparisons reveal emission line excess exclusively in the central region of radio AGN hosts, especially in low mass and radio-compact RDAGN. The high ratio of the excessive [N II] and H supports the scenario that the emission line excess is linked to AGN or shock ionization, but it remains to be understood why the excess is not dependent upon jet luminosity.
@@ Acknowledgements.
- Delvecchio et al. (2021) Delvecchio, I., Daddi, E., Sargent, M. T., et al. 2021, A&A, 647, A123
- Dey et al. (2019) Dey, A., Schlegel, D. J., Lang, D., et al. 2019, AJ, 157, 168
@@
- Harrison & Ramos Almeida (2024) Harrison, C. M. & Ramos Almeida, C. 2024, Galaxies, 12, 17
@@
- Piotrowska et al. (2022) Piotrowska, J. M., Bluck, A. F. L., Maiolino, R., & Peng, Y. 2022, MNRAS, 512, 1052
@@
- Sánchez et al. (2022) Sánchez, S. F., Barrera-Ballesteros, J. K., Lacerda, E., et al. 2022, ApJS, 262, 36
@@ Appendix A Fitting the -SFR relation
To build the -SFR relation, the key is to select a sample of pure star-forming galaxies with reliable SFR measurement and without radio jet contribution.
Using the integral field spectra from MaNGA, the SFR can be derived from the H line, and corrected for dust attenuation on sub-galactic scales. There are galaxies in which the H is not dominated by star formation but instead is possibly contaminated by narrow-line AGN or diffuse ionized gas. We exclude these galaxies using the widely used emission line diagnostic (BPT diagram, Baldwin et al. 1981) based on the [N II]6585/H and [O III]5008/H line ratios measured from the central 2.5″ MaNGA spaxels. Specifically, our ‘pure-SFG’ sample only includes galaxies locating in the star-forming region of the BPT diagram, using the separation diagnostic derived by (Kauffmann et al. 2003), and the galaxies should also have nuclear H equivalent widths larger than 3, which ensures the SFR calculation is not affected by the diffused gas from evolved stars (Cid Fernandes et al. 2011).
Then we visually check the LoTSS image of these galaxies and exclude those showing jet-like features and require the signal-to-noise ratio of radio detection larger than 5. We note that this visual classification cannot exclude unresolved radio AGN. However, they are mostly quiescent galaxies in low redshift (e.g., Fig. 5) and thus will be excluded from the SFG sample by the H equivalent width cut. Considering that there are only very few unresolved radio AGN in the final SFG sample (e.g., some outliers in Fig. 5), the radio AGN contamination for the fitting should be marginal. These unresolved radio AGN could be one source of the intrinsic scatter of the relation and could make the relation slightly overestimated. This way we get a sample of 906 secure SFGs with reliable SFR and radio luminosity measurements.
Fig. 15 shows the emission line selection of this SFG sample. BPT is the distance to the division line in the BPT diagram (Equation 1 in Kauffmann et al. 2003), and the negative values represent that galaxies locate in the star-forming region. As a result of all the constrains, this SFG sample has a SFR limit about 0.1 (see Fig. 2), thus our result about -SFR relation may not be applicable for SFGs with SFRs lower than this value.
We set three parameters in the linear fitting of the -SFR relation. In addition to the slope and constant, we introduce an intrinsic scatter parameter, which should be the scatter level of the -SFR relation after excluding the effect from measurement errors. Our fitting follows the method in Hogg et al. (2010). The Markov chain Monte Carlo (MCMC) sampling result is shown in Fig. 16. The intrinsic error is found to be small, about 0.23 dex, which indicates that the radio emission is a good tracer for galaxy global SFR. The relation is listed in Equation 1. At the 3 level, our result is consistent with previous results based on different SFR tracers and radio data from different frequencies (Wang et al. 2016; Brown et al. 2017; Gürkan et al. 2018; Smith et al. 2021; Heesen et al. 2022; Best et al. 2023).